# Merging LoRAs like Playing LEGO: Pushing the Modularity of LoRA to Extremes Through Rank-Wise Clustering

### v1

@@ 1 Introduction
Previous methods for composing multiple LoRAs have primarily focused on assembling separate LoRAs tailored to specific downstream tasks, which generally require additional training Wu et al. (2023); Wang et al. (2024); Chronopoulou et al. (2023); Yadav et al. (2024a); Huang et al. (2023). Model merging Tang et al. (2024); Yadav et al. (2024b); Ilharco et al. (2022); Yang et al. (2024) offers an alternative approach by aggregating the parameters of multiple LoRAs into a unified adapter without extra training, producing a unified LoRA with more comprehensive capabilities. However, these methods typically employ element-wise parameter fusion, which can neglect and disrupt the internal semantic structure within LoRA. This disruption potentially leads to parameter interference (as discussed in §2.3), thereby hindering the performance of merged LoRA. This paper approaches LoRA merging from a novel perspective, focusing on the fine-grained modularization of LoRA by decomposing it into independent units, which enables the flexible reconstruction of a unified LoRA with comprehensive capabilities.
@@
Figure 1: Further Modularization of LoRA: a) Each LoRA can be further modularized into multiple Minimal Semantic Units (MSUs), each corresponding to a column in ${\bm{A}}$ matrix and a row in matrix ${\bm{B}}$, differentiated by distinct colors. b) The MSUs within a LoRA display permutation invariance, implying that any rearrangement of the MSUs does not affect the output generated by the LoRA. c) Multiple LoRAs exhibit Concatenation-Summation Equivalence, indicating that the summation of outputs from various LoRAs is equivalent to the output of a singular LoRA constructed by concatenating their MSUs.
@@
As illustrated in Fig.1, our motivation for further modularizing LoRA stems from the following insights: a) A rank of LoRA corresponds to a column in the down-projection matrix ${\bm{A}}$ and a row in the up-projection matrix ${\bm{B}}$. Therefore, we consider the parameters associated with each rank as a cohesive entity and define them as Minimal Semantic Units (MSUs), which serve as the fundamental building blocks of each LoRA. b) Within each LoRA, the MSUs exhibit the property of Permutation Invariance, indicating that any permutation of MSUs within a LoRA does not affect the adapter’s output. c) LoRA exhibits the Concatenation-Summation Equivalence property, which states that summing the outputs from multiple LoRAs is equivalent to the output of a single higher-ranked LoRA constructed by concatenating the MSUs of these LoRAs.
@@
In this paper, we introduce a novel method called LoRA-LEGO, which is based on the insight that MSUs act as building blocks that form a LoRA and can be disassembled and reassembled like playing with LEGO. LoRA-LEGO consists of three main steps: (1) Grouping MSUs from candidate LoRAs into a MSU pool; (2) Clustering the MSU pool into $k$ clusters, where $k$ is the target rank of the merged LoRA; (3) Constructing the merged LoRA from the centroids of these clusters, with each centroid representing an MSU, thereby setting the merged LoRA’s rank to $k$. LoRA-LEGO enables the flexible combination of LoRAs with arbitrary ranks by clustering similar MSUs, at the same time effectively resolving parameter interference while merging. This approach allows for targeted rank adjustments in the merged LoRA to preserve task-specific knowledge.
@@ 2.2 Further modularization of LoRA
Minimum Semantic Unit of LoRA. Let ${\bm{A}}$ and ${\bm{B}}$ be matrices in a LoRA module. For each index $i$, define the minimum semantic unit of LoRA as the combined vector ${\bm{s}}_{i}=[{\bm{a}}_{i},{\bm{b}}_{i}]$, where ${\bm{a}}_{i}$ is the $i$-th column of ${\bm{A}}$ and ${\bm{b}}_{i}$ is the $i$-th column of ${\bm{B}}^{T}$ (i.e., the transpose of the $i$-th row of ${\bm{B}}$).
In this context, each column of the down-projection matrix ${\bm{A}}$ and its corresponding row in the up-projection matrix ${\bm{B}}$ are treated as a cohesive unit, defined as a Minimum Semantic Unit (MSU). Each MSU contributes to a rank of the LoRA, encapsulating a distinct semantic fragment of the LoRA’s capacity. Through this definition, LoRAs exhibit the following properties.
@@ 2.3 Problem Formulation and Challenges
Another source of parameter interference stems from knowledge conflict during LoRA merging. As depicted on the right side of Fig.2, knowledge conflict occurs when the merged LoRA lacks sufficient parameter space to encapsulate the comprehensive knowledge. This deficiency forces the merging of task-specific MSUs, resulting in parameter interference. To investigate the impact of knowledge conflict during LoRA merging, we conducted an experiment to demonstrate the performance degradation resulting from merging task-specific MSUs. With a base LoRA trained on the CoLA task, we adapted this LoRA for two new tasks (MNLI and MRPC) by appending an additional MSU to create two separate task-specific LoRAs. Throughout the training process for the new tasks, only the newly introduced MSU for each task was trainable. In this way, the only difference between the LoRAs for MNLI and MRPC was the unique MSU added for each, which encapsulated distinct semantic information tailored to each task. This setup was designed to create two task-specific LoRAs that differed only in one MSU, allowing us to observe parameter interference when merging these task-specific MSUs. The results, depicted in Tab.2.3, demonstrated that averaging the task-specific MSUs from the two LoRAs significantly reduced performance on each task. In contrast, maintaining these task-specific MSUs through concatenation preserved the capabilities specific to each original task. This suggests that ideal merging should maintain task-specific MSUs during LoRA merging to prevent knowledge conflict and effectively resolve parameter interference.
@@ 3.1 LoRA-LEGO framework
The initial stage of merging $p$ LoRAs begins by disassembling each LoRA into various MSUs and grouping all the MSUs from each LoRA together. Let $\{{\bm{A}}_{i},{\bm{B}}_{i}\}_{i=1}^{p}$ represent the LoRA parameters of a layer with rank $r_{i}$. Each LoRA module ${{\bm{A}}_{j},{\bm{B}}_{j}}$ contains $r_{j}$ MSUs, denoted by $\{{\bm{s}}_{j1},{\bm{s}}_{j2},\ldots,{\bm{s}}_{jr_{j}}\}$, where ${\bm{s}}_{jl}=[{\bm{A}}_{jl},{\bm{B}}_{lj}^{T}]$. The MSU pool $\Phi$, which includes MSUs from all the LoRAs to be merged, is constructed as $\Phi=\bigcup_{j=1}^{k}\{{\bm{s}}_{j1},{\bm{s}}_{j2},\ldots,{\bm{s}}_{jr_{j}}\}$.
@@
where ${\bm{a}}_{i}$ and ${\bm{b}}_{i}$ are extracted from each centroid ${\bm{\mu}}_{i}=[{\bm{a}}_{i},{\bm{b}}_{i}]$ as per the MSU definition. The reconstructed LoRA module $\{{\bm{A}}^{\prime},{\bm{B}}^{\prime}\}$ addresses parameter interference by aligning MSUs based on their similarity before merging, achieving a flexible rank that encapsulates comprehensive knowledge across various tasks. An interesting point is that our method sits between model merging, which fuses multiple identical models into the same model, and model ensemble, which takes the average of outputs from different modules, achieving a balance between performance and computational efficiency. We provide a detailed discussion of how our method relates to model merging and model ensemble in the Appendix B.
@@ 3.2 Optimal scale of Merged LoRA
The proof of Theorem 3.1 is detailed in the Appendix D. Overall, to ensure that the LoRA output is correctly scaled, we employ two scaling strategies. First, we reweight the parameters to match the average norms of the parameters within each cluster. Second, we rescale the output of each LoRA to the rank of the merged LoRA, thereby maintaining variance consistency with the original LoRA. These dual scaling strategies enable LoRA-LEGO to deliver enhanced and more robust performance.
@@ 4.1 Multi-task Learning
We evaluate the effectiveness of two scaling strategies for the merged LoRA by varying the number of clusters for LoRA-LEGO, noting that the cluster number corresponds to the rank of the merged LoRA. As illustrated in Fig.7, the original computation of LoRA experiences significant performance degradation with increasing rank of the merged LoRA, primarily due to the expansion of variance associated with the increased rank. Additionally, when the rank of the merged LoRA is relatively low, its performance does not reach its optimum due to the degradation of parameter norms. We also present the performance of each scaling strategy and their combination. Applying parameter reweighing can significantly enhance the performance of the merged LoRA; specifically, the performance of a merged LoRA at rank $1r$ improves by 5%. However, performance degradation still occurs at higher LoRA ranks due to variance expansion. Stabilizing the variance by output reweighting significantly increases performance when the rank is high, although it remains suboptimal due to the decrease of parameter norms. Combining these two scaling strategies yields the best results, demonstrating stable and improved performance across varying ranks of the merged LoRA. After these two scaling strategies are applied, the performance of LoRA-LEGO tends to stabilize; therefore, we use $k=2r$ as the default setting.
@@ 4.2 Mixed-task Evaluation
Adopting the LoraRetriever approach Zhao et al. (2024b), we initially retrieve the top-3 LoRAs based on the sentence embedding similarities between each input sample and a corresponding few-shot samples. Following this, we engage in LoRA composition and evaluate various strategies. This analysis underscores the versatility and superior performance of LoRA-LEGO in handling more complex scenarios.
@@
For all methods, we employ a consistent evaluation pipeline. For each instance in the evaluation set, we initially retrieve the top-3 LoRA, followed by the composition of LoRA. We compared the following LoRA composition methods: (1) Weighted Average, (2) Ensemble, (3) Selection (using the top-1 retrieved LoRA), and (4) Ties-Merging.
@@ 5 Related Work
Many works have discussed how to obtain a comprehensive model through model merging from various perspectives. Some works discuss how to find a set of low-loss paths in the parameter space for model parameter interpolation from the perspective of linear mode connectivity Ainsworth et al. (2022); Entezari et al. (2021). From a similar perspective, we further utilized properties of MSUs, mploying clustering algorithms to provide a flexible solution for enhancing the parameter connectivity during LoRA merging. Additionally, many works attempt to coordinate models trained in a decentralized and separated manner through model merging, utilizing their knowledge transfer capabilities to obtain a model with comprehensive abilities Tang et al. (2024); Don-Yehiya et al. (2022); Yadav et al. (2024b); Matena & Raffel (2022); Jin et al. (2022); Yang et al. (2023a). Recently, with the rise of large language models, more and more works have focused on how to use model aggregation, especially the aggregation of LoRA Chronopoulou et al. (2023); Huang et al. (2023); Zhao et al. (2024b); Wang et al. (2024), to strategically utilize models adapted to multiple domains. These efforts often overlook the parameter interference that occurs during LoRA merging, and some of them require extensive additional training or adaptation. This leads to suboptimal performance in such scenarios or restricts their applicability.
@@
LoRA merging can be applied in various scenarios. For instance, in multi-task learning Tang et al. (2024); Don-Yehiya et al. (2022), models adapt to different domains in a decentralized manner using LoRA, subsequently acquiring multi-task capabilities through merging. In mixed-task scenarios Zhao et al. (2024b; a), LoRAs from diverse domain tasks are uploaded to a centralized service platform, where the service retrieves and composes LoRAs to deliver personalized services based on downstream requests. In federated learning Chen et al. (2023); Zhang et al. (2024), edge devices train LoRAs on private data and upload them to a central server for merging and distribution, enabling iterative optimization through this process. During the alignment phase, RLHF training is conducted for various preferences to obtain multiple LoRAs that meet different requirements. This is followed by providing personalized alignment models through parameter interpolation Jang et al. (2023).
@@ 6 Conclusion
Potential future work includes exploring alternative distance metrics for LoRA-LEGO, such as optimal transport, to provide a more detailed characterization of parameter similarities beyond the typical Euclidean distance. Additionally, further modularization of LoRA could enhance various applications, particularly in federated learning where strategies to minimize communication overhead and expedite model convergence through sharing and aggregating MSUs might be explored. We believe these advancements can significantly benefit a wide range of fields and applications.
@@ Appendix A Difference between LoAR Merging Setting and Model Merging Setting
Previous work on model merging primarily focused on integrating separately trained models to form a comprehensive system. These methods typically involve reloading LoRA parameters into the original model before merging, which introduces additional overhead by necessitating the reconstruction of a corresponding LLM for each LoRA. In many cases, the goal of LoRA merging is to create a new LoRA that consolidates the capabilities of all involved LoRAs for simplified task-specific usage. In contrast, the LoRA merging approach presented in this paper bypasses this step; it directly merges the LoRA parameters to generate a unified LoRA with enhanced capabilities.

### v3

@@ 1 Introduction
Previous methods for composing multiple LoRAs have primarily focused on assembling separate LoRAs tailored to specific downstream tasks, which generally require additional training Wu et al. (2023); Wang et al. (2024); Chronopoulou et al. (2023); Yadav et al. (2024a); Huang et al. (2023). Model merging Tang et al. (2024); Yadav et al. (2024b); Ilharco et al. (2022); Yang et al. (2024) offers an alternative approach by aggregating the parameters of multiple LoRAs into a unified adapter without extra training, producing a unified LoRA with comprehensive capabilities. However, these methods typically employ element-wise parameter fusion, which can neglect and disrupt the internal semantic structure within LoRA. This disruption potentially leads to parameter interference (as discussed in §2.3), thereby hindering the performance of merged LoRA. This paper approaches LoRA merging from a novel perspective, focusing on the fine-grained modularization of LoRA by decomposing it into independent units, which enables the flexible reconstruction of a unified LoRA with comprehensive capabilities.
@@
Figure 1: Further Modularization of LoRA: a) Each LoRA can be further modularized into multiple Minimal Semantic Units (MSUs), each corresponding to a row in ${\bm{A}}$ matrix and a column in matrix ${\bm{B}}$, differentiated by distinct colors. b) The MSUs within a LoRA display permutation invariance, implying that any rearrangement of the MSUs does not affect the output generated by the LoRA. c) Multiple LoRAs exhibit Concatenation-Summation Equivalence, indicating that the summation of outputs from various LoRAs is equivalent to the output of a singular LoRA constructed by concatenating their MSUs.
@@
As illustrated in Fig.1, our motivation for further modularizing LoRA stems from the following insights: a) Each rank in LoRA corresponds to a row in the down-projection matrix $\mathbf{A}$ and a column in the up-projection matrix $\mathbf{B}$. Since the calculations for each rank are independent, we consider the parameters associated with each rank as a cohesive entity. We define these entities as Minimal Semantic Units (MSUs), which serve as the fundamental building blocks of LoRA. b) Within each LoRA, the MSUs exhibit the property of Permutation Invariance, indicating that any permutation of MSUs within a LoRA does not affect the adapter’s output. c) LoRA exhibits the Concatenation-Summation Equivalence property, which states that summing the outputs from multiple LoRAs is equivalent to the output of a single higher-ranked LoRA constructed by concatenating all the MSUs of these LoRAs.
@@
In this paper, we introduce a novel method called LoRA-LEGO, which is based on the insight that MSUs act as building blocks that form a LoRA and can be disassembled and reassembled like playing with LEGO. LoRA-LEGO consists of three main steps: (1) Grouping MSUs from candidate LoRAs into a MSU pool; (2) Clustering the MSU pool into $k$ clusters, where $k$ is the target rank of the merged LoRA; (3) Constructing the merged LoRA from the centroids of these clusters, with each centroid representing an MSU, thereby setting the merged LoRA’s rank to $k$. LoRA-LEGO enables the flexible combination of LoRAs with arbitrary ranks by clustering similar MSUs, at the same time effectively resolving parameter interference while merging. This approach allows for targeted rank adjustments in the merged LoRA to preserve task-specific knowledge. We also observed that variations in parameter norms and the rank size of the merged LoRA affect the output scale. To address this, we implement a dual reweighting strategy that adjusts both the parameters and the outputs, ensuring optimal scaling for the merged LoRA.
@@ 2.2 Further modularization of LoRA
Minimum Semantic Unit of LoRA. Let ${\bm{A}}$ and ${\bm{B}}$ be matrices in a LoRA module. For each index $i$, define the minimum semantic unit of LoRA as the combined vector ${\bm{s}}_{i}=[{\bm{a}}_{i},{\bm{b}}_{i}]$, where ${\bm{a}}_{i}$ is the $i$-th row of ${\bm{A}}$ and ${\bm{b}}_{i}$ is the $i$-th row of ${\bm{B}}^{T}$ (i.e., the transpose of the $i$-th column of ${\bm{B}}$).
In this context, each row of the down-projection matrix ${\bm{A}}$ and its corresponding column in the up-projection matrix ${\bm{B}}$ are treated as a cohesive unit, defined as a Minimum Semantic Unit (MSU). Each MSU contributes to a rank of the LoRA, encapsulating a distinct semantic fragment of the LoRA’s capacity. Through this definition, LoRAs exhibit the following properties.
@@ 2.3 Problem Formulation and Challenges
Another source of parameter interference stems from knowledge conflict during LoRA merging. As depicted on the right side of Fig.2, knowledge conflict occurs when the merged LoRA lacks sufficient parameter space to encapsulate the comprehensive knowledge. This deficiency forces the merging of task-specific MSUs, resulting in parameter interference. To investigate the impact of knowledge conflict during LoRA merging, we conducted an experiment to demonstrate the performance degradation resulting from merging task-specific MSUs. With a base LoRA trained on the CoLA task, we adapted this LoRA for two new tasks (MNLI and MRPC) by appending an additional MSU to create two separate task-specific LoRAs. Throughout the training process for the new tasks, only the newly introduced MSU for each task was trainable. In this way, the only difference between the LoRAs for MNLI and MRPC was the unique MSU added for each, which encapsulated distinct semantic information tailored to each task. This setup was designed to create two task-specific LoRAs that differed only in one MSU, allowing us to observe parameter interference when merging these task-specific MSUs. The results, depicted in Tab.2, demonstrated that averaging the task-specific MSUs from the two LoRAs significantly reduced performance on each task. In contrast, maintaining these task-specific MSUs through concatenation preserved the capabilities specific to each original task. This suggests that ideal merging should maintain task-specific MSUs during LoRA merging to prevent knowledge conflict and effectively resolve parameter interference.
@@ 3.1 LoRA-LEGO framework
The initial stage of merging $p$ LoRAs begins by disassembling each LoRA into various MSUs and grouping all the MSUs from each LoRA together. Let $\{{\bm{A}}_{i},{\bm{B}}_{i}\}_{i=1}^{p}$ represent the LoRA parameters of a layer with rank $r_{i}$. Each LoRA module ${{\bm{A}}_{j},{\bm{B}}_{j}}$ contains $r_{j}$ MSUs, denoted by $\{{\bm{s}}_{j1},{\bm{s}}_{j2},\ldots,{\bm{s}}_{jr_{j}}\}$, where ${\bm{s}}_{jl}=[{\bm{a}}_{jl},{\bm{b}}_{jl}]$ with ${\bm{a}}_{jl}={\bm{A}}_{j}[:,l]$ and ${\bm{b}}_{jl}={\bm{B}}_{j}[l,:]^{T}$. The MSU pool $\Phi$, which includes MSUs from all the LoRAs to be merged, is constructed as $\Phi=\bigcup_{j=1}^{k}\{{\bm{s}}_{j1},{\bm{s}}_{j2},\ldots,{\bm{s}}_{jr_{j}}\}$.
@@
where ${\bm{a}}_{i}$ and ${\bm{b}}_{i}$ are extracted from each centroid ${\bm{\mu}}_{i}=[{\bm{a}}_{i},{\bm{b}}_{i}]$ as per the MSU definition. The reconstructed LoRA module $\{{\bm{A}}^{\prime},{\bm{B}}^{\prime}\}$ addresses parameter interference by aligning MSUs based on their similarity before merging, achieving a flexible rank that encapsulates comprehensive knowledge across various tasks. An interesting point is that our method sits between model merging, which fuses multiple identical models into a singular model, and model ensemble, which takes the average of outputs from different modules, achieving a balance between performance and computational efficiency. We provide a detailed discussion of how our method relates to model merging and model ensemble in the Appendix B.
@@ 3.2 Optimal scale of Merged LoRA
The proof of Theorem 3.1 is detailed in the Appendix D. Overall, to ensure that the LoRA output is correctly scaled, we employ two scaling strategies. First, we reweight the parameters to match the average norms of the parameters within each cluster. Second, we rescale the output of the merged LoRA for maintaining variance consistency with the original LoRA. These dual scaling strategies enable LoRA-LEGO to deliver enhanced and more robust performance.
@@ 4.1 Multi-task Learning
We evaluate the effectiveness of two scaling strategies for the merged LoRA by varying the number of clusters for LoRA-LEGO, noting that the cluster number corresponds to the rank of the merged LoRA. As illustrated in Fig.7, the original computation of LoRA experiences significant performance degradation with increasing rank of the merged LoRA, primarily due to the expansion of variance associated with the increased rank. Additionally, when the rank of the merged LoRA is relatively low, its performance does not reach its optimum due to the degradation of parameter norms. We also present the performance of each scaling strategy and their combination. Applying parameter reweighting can significantly enhance the performance of the merged LoRA when the rank is relatively low; specifically, the performance of a merged LoRA at rank $1r$ improves by 5%. However, as the rank increases, eliminating norm decay more severely exposes variance expansion because norm decay can alleviate this phenomenon, leading to greater performance degradation. Stabilizing the variance by output reweighting significantly increases performance when the rank is high, although it remains suboptimal due to the decrease of parameter norms. Combining these two scaling strategies yields the best results, demonstrating stable and improved performance across varying ranks of the merged LoRA. After these two scaling strategies are applied, the performance of LoRA-LEGO tends to stabilize; therefore, we use $k=2r$ as the default setting.
@@ 4.2 Mixed-task Evaluation
Adopting the LoraRetriever approach Zhao et al. (2024b), we initially retrieve the top-3 LoRAs based on the sentence embedding similarities between each input sample and LoRA’s few-shot samples. Following this, we engage in LoRA composition and evaluate various strategies. This analysis underscores the versatility and superior performance of LoRA-LEGO in handling more complex scenarios.
@@
For all methods, we employ a consistent evaluation pipeline. For each instance in the evaluation set, we initially retrieve the top-3 LoRA, followed by the composition of LoRA. We compared the following LoRA composition methods: (1) Weight Average, (2) Ensemble, (3) Selection (using the top-1 retrieved LoRA), and (4) Ties-Merging.
@@ 5 Related Work
Many works have discussed how to obtain a comprehensive model through model merging from various perspectives. Some works discuss how to find a set of low-loss paths in the parameter space for model parameter interpolation from the perspective of linear mode connectivity Ainsworth et al. (2022); Entezari et al. (2021). From a similar perspective, we further utilized properties of MSUs, employing clustering algorithms to provide a flexible solution for enhancing the parameter connectivity during LoRA merging. Additionally, many works attempt to coordinate models trained in a decentralized and separated manner through model merging, utilizing their knowledge transfer capabilities to obtain a model with comprehensive abilities Tang et al. (2024); Don-Yehiya et al. (2022); Yadav et al. (2024b); Matena & Raffel (2022); Jin et al. (2022); Yang et al. (2023a). Recently, with the rise of large language models, more and more works have focused on how to use model aggregation, especially the aggregation of LoRA Chronopoulou et al. (2023); Huang et al. (2023); Zhao et al. (2024b); Wang et al. (2024), to strategically utilize models adapted to multiple domains. These efforts often overlook the parameter interference that occurs during LoRA merging, and some of them require extensive additional training or adaptation. This leads to suboptimal performance in such scenarios or restricts their applicability.
@@
LoRA merging can be applied in various scenarios. For instance, in multi-task learning Tang et al. (2024); Don-Yehiya et al. (2022), models adapt to different domains in a decentralized manner using LoRA, subsequently acquiring multi-task capabilities through merging. In mixed-task scenarios Zhao et al. (2024b; a), LoRAs from diverse domain tasks are uploaded to a centralized service platform, where the service retrieves and composes LoRAs to deliver personalized services based on downstream requests. In federated learning Chen et al. (2023); Zhang et al. (2024), edge devices train LoRAs on private data and upload them to a central server for merging and distribution, enabling iterative optimization through this process. During the alignment phase, Reinforcement Learning from Human Feedback (RLHF) training is conducted to obtain multiple LoRA models that meet different requirements based on various preferences. Subsequently, personalized alignment models can be provided through parameter interpolation, as discussed in  Jang et al. (2023).
@@ 6 Conclusion
Potential future work includes exploring alternative distance metrics for LoRA-LEGO, such as optimal transport, to better characterize parameter similarities beyond the standard Euclidean distance. Additionally, further modularization of LoRA could enhance various applications. For example, in federated learning, strategies to minimize communication overhead and expedite model convergence through sharing and aggregating MSUs could be explored. We believe these advancements could significantly benefit a wide range of fields and applications.
@@ Appendix A Difference between LoAR Merging Setting and Model Merging Setting
Previous work on model merging primarily focused on integrating separately trained models to form a comprehensive system. These methods typically involve reloading LoRA parameters into the original model before merging, which introduces additional overhead by necessitating the reconstruction of a corresponding LLM for each LoRA. In many cases, the goal of LoRA merging is to create a new LoRA that consolidates the capabilities of all involved LoRAs for simplified task-specific usage. In contrast, the LoRA merging setting presented in this paper bypasses the LoRA reload step; it directly merges the LoRA parameters to construct a unified LoRA with comprehensive capabilities.